OpenStudy (anonymous):

order each quadratic function from widest to narrowest graph. 10. y=3x^2, y= 1/2x^2, y=4x^2

OpenStudy (anonymous):

\[\left\{y=\frac{x^2}{2},y=3 x^2,y=4 x^2\right\} \]Plot attached. The color for the answers from left to right is red, blue, yellow.

OpenStudy (anonymous):

Plot is attached to this answer box.